OpenStudy (31356):

(x-8)(x-2) Multiplying polynomials MEDAL REWARDED!!

OpenStudy (31356):

@Abdulhameed

OpenStudy (solomonzelman):

For example (y+4)(y+2)= y(y+2) + 4(y+2) y^2 +2y + 4y + 8 add like terms (with the y 's power of 1) y^2 + 6y +8 See?

OpenStudy (31356):

Ok

OpenStudy (solomonzelman):

Try to do your problem the same way, I'll start you off. (x -8) (x-2) = x * (x-2) + 8 * (x-2) = x^2 - 2x + (-8x) + (-16) can you continue?

OpenStudy (31356):

Yes

OpenStudy (solomonzelman):

Tell me what you get at the end, to make sure you are correct.